# Higher-order Gini indices: An axiomatic approach

### v1

@@ Abstract
Via an axiomatic approach, we characterize the family of $n$-th order Gini deviation, defined as the expected range over $n$ independent draws from a distribution, to quantify joint dispersion across multiple observations. This extends the classical Gini deviation, which relies solely on pairwise comparisons. Our generalization grows increasingly sensitive to tail inequality as $n$ increases, offering a more nuanced view of distributional extremes. We show that these higher-order Gini deviations admit a Choquet integral representation, inheriting the desirable properties of coherent deviation measures. Furthermore, we prove that both the $n$-th order Gini deviation and its normalized version, the $n$-th order Gini coefficient, are $n$-observation elicitable, facilitating rigorous backtesting. Empirical analysis using World Inequality Database data reveals that higher-order Gini coefficients detect disparities obscured by the classical Gini coefficient, particularly in cases of extreme income or wealth concentration. Our results establish higher-order Gini indices as valuable complementary tools for robust inequality assessment.
@@ 1 Introduction
To better understand these statements, we plot in Figure 2 the wealth shares held by the top 1% and top 10% of the population in China, Canada, and UK. The data clearly show that both the top 1% and top 10% populations in China have consistently held a larger share of total wealth than their counterparts in Canada and the UK after 2010. This provides strong evidence against the first two statements, indicating that China’s wealth distribution has been notably more unequal than UK and Canada. On the other hand, both Figures 1 and 2 agree that the wealth distribution in China may be stabilized after 2010.
@@
where $X_{1},\dots,X_{n}$ are iid copies of XX. [^3] During the writing of our paper, this formula also appeared in Gavilan-Ruiz et al. (2024), with minimal technical overlap; see Remark 3 for a detailed discussion. The proof of this axiomatic characterization goes by first establishing a signed Choquet integral representation and then identifying those that can be represented by expected values of functions of iid observations. Many classes of decision models, such as the dual utility of Yarri (1987) and the Choquet expected utility model of Schmeidler (1989), are formulated based on signed Choquet integrals; see the more recent works of Strack and Wambach (2017), Gershkov et al. (2025) and Hartmann and Kauffeldt (2025). The main differences here is that our functionals $\mathrm{GD}_{n}$ do not satisfy monotonicity, thus they are called “signed”.
@@
Many variants of GD and GC exist in the literature; see Ceriani and Verme (2012) for several formulations of GC proposed by Gini, and Yitzhaki (1998, 2003), Yitzhaki and Schechtman (2012) and Vila and Saulo (2025) for further developments. Our generalizations of GD and GC provide a unique perspective by altering the number of observations used in its calculation, while keeping all nice properties of GD and GC.
After reviewing the classical Gini deviation and Gini coefficient in Section 2, we first present the axioms satisfied by the Gini indices and then develop a rigorous axiomatic framework for their higher-order generalizations in Section 3. In Section 4, we explore the key properties of the higher-order Gini indices in detail. First, we examine the economic properties of the higher-order Gini coefficient, emphasizing its interpretability and practical relevance. Second, we show that $\mathrm{GD}_{n}$ admits a representation as a signed Choquet integral with a concave distortion function. Moreover, $\mathrm{GD}_{n}$ belongs to the class of deviation measures in the sense of Rockafellar et al. (2006) and the class of coherent measures of variability in the sense of Furman et al. (2017).
This integral formulation guarantees several desirable properties, including symmetry, convexity, and comonotonic additivity, and leads to economically useful properties of $\mathrm{GC}_{n}$, such as quasi-convexity and mixture-quasi-concavity. Finally, we present a comparative analysis between the $n$-th order GD and the standard deviation.
@@ 3.1 Axioms of the Gini indices
Note that [A12] is made so that $\rho/\mathbb{E}$ can be readily used as the corresponding relative index playing the role of GC. Therefore, the Gini axioms are made for the pair $(\rho,\rho/\mathbb{E})$ in place of $(\mathrm{GD},\mathrm{GC})$, although the mathematical statements are written for $\rho$ for simplicity. The economic interpretation of some of these axioms will be discussed in Section 4.
@@ 3.2 Axiomatic characterization
and therefore, $\mathrm{GD}_{2}=\mathrm{GD}$. Following this, we define the $n$-th order Gini coefficient $\mathrm{GC}_{n}$ for $n\geq 2$ by
@@
Note that $\mathrm{GC}_{n}$ has a natural range of $[0,1)$, as shown in Theorem 1, which relies on the following example. This example also justifies that $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ bear the same interpretation as $\mathrm{GD}$ and $\mathrm{GC}$ when it comes to measuring wealth or income inequality, as the worst-case scenario is attained asymptotically by the distribution in which a tiny ($\epsilon$ in the example) proportion of the population has all the wealth or income.
@@
In Theorem 1, we have shown that if $(a_{1},\dots,a_{n})$ is in $\Delta_{n}$, then all the Gini axioms are satisfied. One may wonder whether the reverse holds true, that is, whether for $\rho$ in (3), [A1]–[A12] jointly imply $(a_{1},\dots,a_{n})\in\Delta_{n}$. It turns out that this is not true, even removing the redundant terms (e.g., $\mathrm{GD}_{1}=\mathrm{GD}_{2}=\mathrm{GD}_{3}$). For an example, define $\rho(X)=2{\rm GD}_{2}(X)-{\rm GD}_{4}(X)$. By Theorem 3, we know that $\rho$ is a signed Choquet integral whose distortion function is given by
@@ 4.2 Representation as signed Choquet integrals
In the next result, we represent $\mathrm{GD}_{n}$ as either a signed Choquet integral or a covariance. This is useful in proving several properties of $\mathrm{GD}_{n}$ in Theorem 1. In what follows, let $U_{X}$ be a uniform random variable such that $F_{X}^{-1}(U_{X})=X$ almost surely, and its existence is guaranteed by Föllmer and Schied (2016, Lemma A.32).
Theorem 3.
@@ 5 Elicitability and multi-observation elicitability
When $n$ is an odd number, we can check that the $n$-th degree term $t^{n}$ in the distortion function $h_{n}$ in (5) get canceled. Therefore, in this case, by using Theorem 5, $\mathrm{GC}_{n}$ is (n1)(n-1)-observation elicitable, which is a stronger condition than $n$-observation elicitability, as shown in Proposition 3. Note that the score function for n1n-1 observations is different from the one obtained in Theorem 4 for $n$ observations.
@@ 8 Conclusion
Motivated by the representation of the classical Gini deviation as the expectation of a function of two independent copies of a random variable, this work presents a systematic generalization—developed through an axiomatic framework—of the classical Gini deviation and Gini coefficient to their higher-order counterparts, thereby enriching the analytical toolkit for inequality and dispersion measurement. Specifically, we consider natural extensions involving expectations of functions of $n\geq 2$ independent and identically distributed copies. By imposing the properties of symmetry, comonotonic additivity, and continuity, we provide an axiomatic characterization showing that any such measure can be represented as an affine combination of the higher-order Gini deviations. By transitioning from pairwise to groupwise comparison structures, the proposed indices capture subtle aspects of tail concentration and extremal behavior—features often obscured under traditional two-observation paradigms.
We analyze the economic properties of the higher-order Gini coefficient, and show that higher-order Gini deviations admit Choquet integral representations, satisfying key properties of coherent deviation and risk measures. The relationship between the high order Gini deviation and the standard deviation is explored. Furthermore, we establish their $n$-observation elicitability, enabling formal comparative backtesting and bridging the gap between descriptive indices and decision-theoretic evaluation. Empirical analyses using global income and wealth data illustrate the strength of this framework: while the classical GC often masks nuanced disparities across regions or subpopulations, higher-order indices reveal meaningful differences—particularly among top earners. These findings highlight the value of $\mathrm{GC}_{n}$ as a diagnostic tool for policymakers, regulators, and data scientists facing increasingly complex distributional challenges.
As data environments grow more intricate, the flexibility and theoretical depth of higher-order Gini indices make them promising instruments for capturing refined patterns of inequality, concentration, and systemic risk. In risk management, these indices provide coherent, tail-sensitive alternatives to variance-based risk measures. Their emphasis on upper-tail risks aligns naturally with the objectives of solvency analysis, reinsurance pricing, and regulatory capital assessment. Meanwhile, in machine learning, the multi-observation structure supports the design of loss functions that better capture group-level dispersion, promoting robustness and fairness in algorithmic decision-making—especially in settings involving skewed or heavy-tailed data.
Remark 3.
After completing the first draft of this paper, we became aware of the study of Gavilan-Ruiz et al. (2024), which contains the same definition of the higher-order Gini indices. Our work was developed independently, with the first draft, containing the key concepts, written and privately communicated in September 2022. The technical overlap between our paper and Gavilan-Ruiz et al. (2024) is minimal. Gavilan-Ruiz et al. (2024) obtained monotonicity of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ with respect to the order parameter $n$, which we reported in Appendix C for completeness. All results presented in the main text of our paper are new to the literature.
@@ Appendix A Proofs
A.1 Proof of Theorem 1
We first prove the equivalence statement. Some statements rely on properties of $\mathrm{GD}_{n}$ in Theorem 3 in Section 4, which will be proved later separately.
@@ A.1 Proof of Theorem 1
This implies that $\rho$ satisfies [A1]. It follows from Theorem 3 that $\rho$ defined by (3) is a signed Choquet integral, and thus, it follows from Theorem 1 of Wang et al. (2020a) that $\rho$ satisfies symmetry, comonotonic additivity, and uniform norm continuity. This completes the proof of sufficiency.
@@
By the representation result of ${\rm{GD}}_{i}$ in terms of signed Choquet integral in Theorem 3 of Section 4.2, $\rho$ admits the form of (3) with $a_{1}+a_{2}=-c_{2}$ and $a_{i}=-ic_{i}/2$ for $i=3,\dots,n$.
Sufficiency for other axioms. Nonnegativity [A5] follows from (4) and $a_{1},\dots,a_{n}\geq 0$. [A6]–[A7] follow directly from properties of signed Choquet integrals in Wang et al. (2020a). [A8]–[A11] follow the fact that each $\mathrm{GD}_{i}$ has a concave distortion function in Theorem 3 and $a_{1},\dots,a_{n}\geq 0$; see Theorem 3 of Wang et al. (2020a). To show [A12], note that $\mathbb{E}[\max\{X_{1},\dots,X_{n}\}]<\mathbb{E}[X_{1}+\dots+X_{n}]=n\mathbb{E}[X_{1}]$, and hence $\mathrm{GD}_{i}(X)/\mathbb{E}[X]\leq 1$ for all $X\in\mathcal{X}_{+}$. Example 1 below illustrates that the functionals $\mathrm{GD}_{i}(X)/\mathbb{E}[X]$ share the same sequence $(X_{\epsilon})$ of random variables such that $\mathrm{GD}_{i}(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$ as $\epsilon\downarrow 0$. Therefore, as their convex combination, $\rho$ satisfies $\rho(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$, and hence [A12] holds.
A.2 Proofs of Section 4
@@
(ii) It is clear from (12) that $\mathrm{GD}_{n}$ is a signed Choquet integral with a concave distortion function
$h_{n}(t)=\frac{1}{n}\left(1-t^{n}-(1-t)^{n}\right),~~~~t\in[0,1].$
(iii) Let $U_{X}$ be a uniform random variable such that $F_{X}^{-1}(U_{X})=X$ a.s. Then
$\mathrm{GD}_{n}(X)=\int_{0}^{1}F^{-1}_{X}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t=\mathbb{E}[X(U_{X}^{n-1}-(1-U_{X})^{n-1})].$
Also, note that $\mathbb{E}[U_{X}^{n-1}-(1-U_{X})^{n-1}]=0$, and thus we have
$\mathrm{GD}_{n}(X)=\mathrm{Cov}(X,U_{X}^{n-1}-(1-U_{X})^{n-1}).$
This completes the proof. ∎
@@
By Proposition 5 in Appendix C, since $n\leq m$, we already know $\mathrm{GD}_{n}(X)/\mathrm{GD}_{m}(X)\leq 1$ for $X\in L^{1}_{*}$. For the lower bound, (7) gives, with $h_{n}$ defined in (5),
@@ A.4 Proof of Theorem 6
$\mathrm{Var}\left[\frac{1}{\mathbb{E}[X]}\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}\left(s^{n-1}-(1-s)^{n-1}-\frac{\mathrm{GD}_{n}(X)}{\mathbb{E}[X]}\right)\mathrm{d}s\right]$
@@
Thus, $\sqrt{N}(\widehat{\mathrm{GC}}_{n}(N)-\mathrm{GC}_{n}(X))\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}(0,\sigma^{2}_{\mathrm{GC}_{n}})$ in which $\sigma^{2}_{\mathrm{GC}_{n}}$ is given by (10).

### v2

@@ Abstract
Via an axiomatic approach, we characterize the family of $n$-th order Gini deviation, defined as the expected range over $n$ independent draws from a distribution, to quantify joint dispersion across multiple observations. This family extends the classical Gini deviation, which relies solely on pairwise comparisons. The normalized version is called a high-order Gini coefficient. The generalized indices grow increasingly sensitive to tail inequality as $n$ increases, offering a more nuanced view of distributional extremes. The higher-order Gini deviations admit a Choquet integral representation, inheriting the desirable properties of coherent deviation measures. Furthermore, we show that both the $n$-th order Gini deviation and the $n$-th order Gini coefficient are statistically $n$-observation elicitable, allowing for direct computation through empirical risk minimization. Data analysis using World Inequality Database data reveals that higher-order Gini coefficients capture disparities that the classical Gini coefficient may fail to reflect, particularly in cases of extreme income or wealth concentration.
@@ 1 Introduction
To better understand these statements, we plot in Figure 2 the wealth shares held by the top 1% and top 10% of the population in China, Canada, and UK. The data clearly show that both the top 1% and top 10% populations in China have consistently held a larger share of total wealth than their counterparts in Canada and the UK after 2010. This provides evidence against the first two statements, indicating that China’s wealth distribution has been notably more unequal than UK and Canada. On the other hand, both Figures 1 and 2 agree that the wealth distribution in China may be stabilized after 2010.
@@
where $X_{1},\dots,X_{n}$ are iid copies of $X$. The proof of our axiomatic characterization goes by first establishing a signed Choquet integral representation and then identifying those that can be represented by expected values of functions of iid observations. Many classes of decision models, such as the dual utility of Yaari (1987) and the Choquet expected utility model of Schmeidler (1989), are formulated based on Choquet integrals; see the more recent works of Strack and Wambach (2017), Gershkov et al. (2025) and Hartmann and Kauffeldt (2025). The main differences here is that our functionals $\mathrm{GD}_{n}$ do not satisfy monotonicity, thus they are called “signed”.
@@
The quantity $\mathrm{GC}_{n}$ as a generalization of GC was recently proposed by Gavilan-Ruiz et al. (2024), but a comprehensive study is still lacking, particularly with respect to an axiomatic framework and empirical analysis.
@@
Many other variants of GD and GC exist in the literature; see Ceriani and Verme (2012) for several formulations of GC proposed by Gini, and Yitzhaki (1998, 2003), Yitzhaki and Schechtman (2012) and Vila and Saulo (2025) for further developments.
After reviewing the classical Gini deviation and Gini coefficient in Section 2, we first present the axioms satisfied by the Gini indices and then develop a rigorous axiomatic framework for their higher-order generalizations in Section 3. In Section 4, we explore the key properties of the higher-order Gini indices in detail. First, we examine the economic properties of the higher-order Gini coefficient, emphasizing its interpretability and practical relevance. Second, we show that $\mathrm{GD}_{n}$ admits a representation as a signed Choquet integral with a concave distortion function. Moreover, $\mathrm{GD}_{n}$ belongs to the class of deviation measures in the sense of Rockafellar et al. (2006) and the class of coherent measures of variability in the sense of Furman et al. (2017). This integral formulation guarantees several desirable properties, including symmetry, convexity, and comonotonic additivity, and leads to economically useful properties of $\mathrm{GC}_{n}$, such as quasi-convexity and mixture-quasi-concavity. Finally, we present a comparative analysis between the $n$-th order GD and the standard deviation.
@@ 3.1 Axioms of the Gini indices
Note that [A12] is made so that $\rho/\mathbb{E}$ can be readily used as the corresponding relative index playing the role of GC. Therefore, the Gini axioms are made for the pair $(\rho,\rho/\mathbb{E})$ in place of $(\mathrm{GD},\mathrm{GC})$, although the mathematical statements for $\rho$ are simpler to present, except for [A12]. The economic interpretation of some of these axioms will be discussed in Section 4.
@@ 3.2 Axiomatic characterization
and therefore, $\mathrm{GD}_{2}=\mathrm{GD}$. Following this, the $n$-th order Gini coefficient $\mathrm{GC}_{n}$ for $n\geq 2$, is defined by, as in Gavilan-Ruiz et al. (2024),
@@
Note that $\mathrm{GC}_{n}$ has a natural range of $[0,1)$ due to [A12], and this fact has been observed by Gavilan-Ruiz et al. (2024). The following example explains the extreme distributions, and also justifies that $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ bear the same interpretation as $\mathrm{GD}$ and $\mathrm{GC}$ when it comes to measuring wealth or income inequality, as the worst-case scenario is attained asymptotically by the distribution in which a tiny ($\epsilon$ in the example) proportion of the population has all the wealth or income.
@@
In Theorem 1, we have shown that if $(a_{1},\dots,a_{n})$ is in $\Delta_{n}$, then all the Gini axioms are satisfied. One may wonder whether the reverse holds true, that is, whether for $\rho$ in (3), [A1]–[A12] jointly imply $(a_{1},\dots,a_{n})\in\Delta_{n}$. It turns out that this is not true, even removing the redundant terms (e.g., $\mathrm{GD}_{1}=\mathrm{GD}_{2}=\mathrm{GD}_{3}$). For an example, define $\rho(X)=2{\rm GD}_{2}(X)-{\rm GD}_{4}(X)$. By Proposition 1, we know that $\rho$ is a signed Choquet integral whose distortion function is given by
@@ 4.2 Representation as signed Choquet integrals
In the next proposition, we represent $\mathrm{GD}_{n}$ as a signed Choquet integral. This is useful in proving several properties of $\mathrm{GD}_{n}$ in Theorem 1. In what follows, let $U_{X}$ be a uniform random variable such that $F_{X}^{-1}(U_{X})=X$ almost surely, and its existence is guaranteed by Föllmer and Schied (2016, Lemma A.32).
Proposition 1.
@@
Remark 3.
From (5), we can easily generalize the definition of $\mathrm{GD}_{n}$ to real numbers $n>1$, and these indices are also signed Choquet integrals. Similarly, one can define $\mathrm{GC}_{n}$ for all real numbers $n>1$. Nevertheless, for a non-integer $n$, sample representability [A1] is no longer satisfied by $\mathrm{GD}_{n}$ in (5). For this reason, we focus integer values of $n$.
@@ 5 Elicitability and multi-observation elicitability
When $n$ is an odd number, we can check that the $n$-th degree term $t^{n}$ in the distortion function $h_{n}$ in (6) get canceled. Therefore, in this case, by using Theorem 4, $\mathrm{GD}_{n}$ is (n1)(n-1)-observation elicitable, which is a stronger condition than $n$-observation elicitability, as shown in Proposition 4. Note that the score function for n1n-1 observations is different from the one obtained in Theorem 3 for $n$ observations.
@@ 8 Conclusion
We provided an axiomatic study for a generalization of the classical Gini deviation and Gini coefficientas tools for inequality and dispersion measurement. We considered an axiom involving expectations of functions of $n\geq 2$ independent and identically distributed copies. By imposing the properties of symmetry, comonotonic additivity, and continuity, we obtained an axiomatic characterization showing that any such measure can be represented as an affine combination of the higher-order Gini deviations. By transitioning from pairwise to groupwise comparison structures, the generalized indices capture aspects of tail concentration and extremal behavior, different from the traditional two-observation formulation.
We analyze the economic properties of the higher-order Gini coefficient, and show that higher-order Gini deviations admit Choquet integral representations, satisfying key properties of coherent deviation and risk measures. The relationship between the high order Gini deviation and the standard deviation is explored. Furthermore, we establish their $n$-observation elicitability, bridging the gap between descriptive indices and decision-theoretic evaluation. Empirical analyses using global income and wealth data illustrate the strength of this framework: while the classical GC often masks nuanced disparities across regions or subpopulations, higher-order indices reveal meaningful differences—particularly among top earners. These findings highlight the value of $\mathrm{GC}_{n}$ as a diagnostic tool for policymakers, regulators, and data scientists facing increasingly complex distributional challenges.
As data environments grow more intricate, the flexibility of higher-order Gini indices make them promising instruments for capturing refined patterns of inequality, concentration, and systemic risk. In risk management, these indices provide coherent, tail-sensitive alternatives to variance-based risk measures. Their emphasis on upper-tail risks aligns naturally with the objectives of solvency analysis, reinsurance pricing, and regulatory capital assessment. Meanwhile, in machine learning, the multi-observation structure supports the design of loss functions that better capture group-level dispersion, promoting robustness and fairness in algorithmic decision-making—especially in settings involving skewed or heavy-tailed data.
@@ Appendix A Proofs
Proof of Theorem 1.
We first prove the equivalence statement. Some statements rely on properties of $\mathrm{GD}_{n}$ in Proposition 1 in Section 4, which will be proved later separately.
@@
This implies that $\rho$ satisfies [A1]. It follows from Proposition 1 that $\rho$ defined by (3) is a signed Choquet integral, and thus, it follows from Theorem 1 of Wang et al. (2020a) that $\rho$ satisfies symmetry, comonotonic additivity, and uniform norm continuity. This completes the proof of sufficiency.
@@
By the representation result of ${\rm{GD}}_{i}$ in terms of signed Choquet integral in Proposition 1 of Section 4.2, $\rho$ admits the form of (3) with $a_{1}+a_{2}=-c_{2}$ and $a_{i}=-ic_{i}/2$ for $i=3,\dots,n$.
Sufficiency for other axioms. Nonnegativity [A5] follows from (4) and $a_{1},\dots,a_{n}\geq 0$. [A6]–[A7] follow directly from properties of signed Choquet integrals in Wang et al. (2020a). [A8]–[A11] follow the fact that each $\mathrm{GD}_{i}$ has a concave distortion function in Proposition 1 and $a_{1},\dots,a_{n}\geq 0$; see Theorem 3 of Wang et al. (2020a). To show [A12], note that $\mathbb{E}[\max\{X_{1},\dots,X_{n}\}]<\mathbb{E}[X_{1}+\dots+X_{n}]=n\mathbb{E}[X_{1}]$, and hence $\mathrm{GD}_{i}(X)/\mathbb{E}[X]\leq 1$ for all $X\in\mathcal{X}_{+}$. Example 1 below illustrates that the functionals $\mathrm{GD}_{i}(X)/\mathbb{E}[X]$ share the same sequence $(X_{\epsilon})$ of random variables such that $\mathrm{GD}_{i}(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$ as $\epsilon\downarrow 0$. Therefore, as their convex combination, $\rho$ satisfies $\rho(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$, and hence [A12] holds. ∎
@@
(ii) It follows from (13) that $\mathrm{GD}_{n}$ is a signed Choquet integral with a concave distortion function in (6). ∎
@@
By Proposition 6 in Appendix C, since $n\leq m$, we know $\mathrm{GD}_{n}(X)/\mathrm{GD}_{m}(X)\leq 1$ for $X\in L^{1}_{*}$. For the lower bound, (8) gives, with $h_{n}$ defined in (6),
@@
$\mathrm{Var}\left[\frac{1}{\mathbb{E}[X]}\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}\left(s^{n-1}-(1-s)^{n-1}-\mathrm{GC}_{n}(X)\right)\mathrm{d}s\right]$
@@
Thus, $\sqrt{N}(\widehat{\mathrm{GC}}_{n}(N)-\mathrm{GC}_{n}(X))\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}(0,\sigma^{2}_{\mathrm{GC}_{n}})$ in which $\sigma^{2}_{\mathrm{GC}_{n}}$ is given by (11). ∎